A glass-wedge with the angle of wedge of 30 seconds of an arc is formed With a liquid of

R.1. 1.47. Find the number of dark interterence fringes per cm of the wedge length.
1
Expert's answer
2019-11-07T09:10:55-0500

We can write the condition for minimum interference as

"2hn+\frac {λ}{2}=(2m+1)×\frac {λ}{2} (1)"

where h is the thickness of the wedge, m is the number of dark interterence fringes


Using (1) we got

"2hn=m\u03bb (2)"

θ is the angle between the plates in radians (this angle is small, so \tan {θ} = θ in radians).


In this case, we got

"θ=\frac {h}{l} (3)"


Using (3) we got

"h= \u03b8\u00d7l (4)"

We put (4) in (2)

"2\u03b8ln=m\u03bb (5)"

We got

"\frac {m}{l}=\frac {2θn}{λ} (6)"

In our case, we have θ=1.46×10-4, n=1.47, λ=582 nm


The number of dark interterence fringes per cm of the wedge length is equal to

"\frac {m}{l}=3.68 \frac {fringes}{cm}"